# Multi-frequency analysis of the ALMA and VLA high resolution continuum observations of the substructured disc around CI Tau

### v1

@@
11institutetext: Max Planck Institute for Astronomy, Königstuhl 17, 69117 Heidelberg, Germany 22institutetext: Institute of Astronomy, University of Cambridge, Madingley Road, Cambridge CB3 OHA, UK 33institutetext: Università degli Studi di Milano, Via Giovanni Celoria 16, 20133 Milano, Italy 44institutetext: European Southern Observatory, Karl-Schwarzschild-Str. 2, 85748 Garching bei München, Germany 55institutetext: Departamento de Astronomía, Universidad de Chile, Camino El Observatorio 1515, Las Condes, Santiago, Chile 66institutetext: Institute for Astronomy, University of Hawaii, Honolulu, HI 96822, USA 77institutetext: INAF – Osservatorio Astrofisico di Arcetri, L.go E. Fermi 5, 50125 Firenze, Italy 88institutetext: Astrophysics Group, Department of Physics, Imperial College London, Prince Consort Rd, London SW7 2A2, UK 99institutetext: School of Physics and Astronomy, University of Leeds, Leeds, LS2 9JT, UK 1010institutetext: Leiden Observatory, Leiden University, P.O. Box 9513, 2300 RA Leiden, The Netherlands $1$institutetext: School of Physics & Astronomy, University of Leicester, University Road, Leicester LE1 7RH, UK 1212institutetext: Alma Mater Studiorum Università di Bologna, Dipartimento di Fisica e Astronomia (DIFA), Via Gobetti 93/2, 40129 Bologna, Italy
@@
Preference for sub-mm-sized low-porosity amorphous carbonaceous grains
@@
We present high angular resolution ($50\,{\rm mas}$) and sensitivity Atacama Large Millimeter/submillimeter Array (ALMA) Band 3 ($3.1\,{\rm mm}$) and Very Large Array (VLA) Ka Band ($9.1\,{\rm mm}$) observations of the multi-ringed disc around the $3\,{\rm Myr}$-old solar-mass star CI Tau. These new data were combined with similar-resolution archival ALMA Band 7 ($0.9\,{\rm mm}$) and 6 ($1.3\,{\rm mm}$) observations, and new and archival VLA Q ($7.1\,{\rm mm}$), Ku ($2.0\,{\rm cm}$), X ($3.0\,{\rm cm}$), and C Band ($6.0\,{\rm cm}$) photometry to study the properties of dust in this system. At wavelengths $\leq 3.1\,{\rm mm}$ the continuum emission from CI Tau is very extended ($\geq 200\,{\rm au}$) and highly substructured (with three gaps, four rings, and two additional gap-ring pairs identified by non-parametric visibility modelling). Instead, the VLA Ka Band data are dominated by a centrally peaked bright component, only partially ($\leq 50\%$) due to dust emission, surrounded by a marginally-detected, faint and smooth halo. We fitted the ALMA and VLA Ka Band data together, adopting a physical model that accounts for the effects of dust absorption and scattering. For our fiducial dust composition (“Ricci” opacities), we retrieved a flat maximum grain size distribution across the disc radius, with $a_{\rm max}=(7.1\pm 0.8)\times 10^{-2}\,{\rm cm}$, that we tentatively attributed to fragmentation of fragile dust or bouncing. We tested, for the first time, the dependence of our results on the adopted dust composition model to assess which mixture can best reproduce the observations. We found that the “Ricci” opacities work better than the traditionally adopted “DSHARP” ones, while graphite-rich mixtures perform significantly worse. We also show that, for our fiducial composition, the data prefer low porosity ($\leq 70\%$) grains. This is in contrast with recent claims of highly porous aggregates in younger sources, which we tentatively justified by time-dependent compaction at the fragmentation or bouncing barrier. Our results on composition and porosity are in line with constraints from disc population synthesis models and naturally arise from CI Tau’s peculiar spectral behaviour (i.e., the abrupt steepening of its spectral index at wavelengths longer than $3.1\,{\rm mm}$), making this disc a unique target to characterise the properties of disc solids, thus ideal for deeper cm-wavelength observations and follow-up dust polarisation studies.
@@ 1 Introduction
Dust is an essential ingredient of every planet-formation model. First and foremost, it is the material planet(esimal)s are made of. By dominating the disc opacity, solids also set the temperature structure, and thus the excitation conditions and reaction rates of volatiles (e.g., Gavino et al., 2021; Öberg et al., 2023). In addition, electron/ion recombination and chemical reactions on the surface of grains affect the disc ionisation structure (e.g., Desch & Turner, 2015) and prompt the formation of complex organic molecules (e.g., Öberg & Bergin, 2021). Dust collisional evolution (growth/fragmentation) and transport (vertical settling and radial drift) not only determine the availability of solids to form planetary cores (Brauer et al., 2008; Birnstiel et al., 2010), but also affect the dust-to-gas thermal coupling (Facchini et al., 2017) and the abundance of gas-phase chemical species: as dust migrates radially, the volatiles frozen onto grains undergo thermal desorption upon transition across their snowlines (e.g., Booth et al., 2017; Booth & Ilee, 2019; Krijt et al., 2020; Van Clepper et al., 2022), thus locally enriching the feedstock of exoplanet atmospheres (Madhusudhan, 2019). Therefore, unsurprisingly, determining the physical and chemical properties of protoplanetary disc solids is an essential first step to understand planet formation (e.g., Miotello et al., 2023; Birnstiel, 2024). For this reason, attempts have been made to infer the density, size, distribution, and temperature of disc solids by modelling the continuum emission from planet-forming discs at multiple wavelengths.
@@
Early works, either focusing on a handful of targets of interest, or surveying tens of objects in nearby star-formation regions, measured integrated spectral indices, first in the (sub-)mm, then at progressively longer wavelengths, significantly smaller than in the interstellar medium (ISM). While, especially at (sub-)mm wavelengths, the contribution of optically thick regions could not be ruled out (Ricci et al., 2012), these results were generally interpreted (Beckwith & Sargent, 1991; Andrews & Williams, 2005; Rodmann et al., 2006), sometimes in combination with information on the (wavelength dependence of the) disc radius (e.g., Testi et al., 2003; Wilner et al., 2005; Ricci et al., 2010a, b; Isella et al., 2010; Banzatti et al., 2011; Guilloteau et al., 2011), as a consequence of substantial grain growth up to mm/cm sizes, depending on the adopted dust optical properties (see Testi et al. 2014 and Andrews 2020 for a summary picture). In addition, the first multi-wavelength studies able to resolve few bright sources from (sub-)mm to cm wavelengths (Pérez et al., 2012, 2015; Trotta et al., 2013; Tazzari et al., 2016) inferred radially decreasing grain size profiles, in qualitative agreement with the predictions of dust evolution models (Brauer et al., 2008; Birnstiel et al., 2010).
@@
The advent of the Atacama Large Millimeter/submillimeter Array (ALMA) substantially revolutionised this picture. One of ALMA’s major breakthroughs was the almost ubiquitous detection of small scale structures in the continuum emission of (the largest/brightest) protoplanetary discs (Andrews, 2020; Bae et al., 2023). At (sub-)mm wavelengths, substructures are most commonly detected in the form of axisymmetric troughs and peaks (colloquially referred to as “gaps and rings”, e.g., Andrews et al. 2016; Long et al. 2018; Andrews et al. 2018; Huang et al. 2018b) or cavities (e.g., Keppler et al., 2019; Facchini et al., 2020), but spirals (Pérez et al., 2016; Huang et al., 2018c; Kurtovic et al., 2018; Paneque-Carreño et al., 2021) and crescents (e.g., van der Marel et al., 2013; Isella et al., 2018; Pérez et al., 2018; Dong et al., 2018) are sometimes also present. On the one hand, the most tantalising interpretation of these structures is that they trace dynamical interactions between massive (0.1 to $10.0\,M_{\rm Jup}$, Lodato et al. 2019; Bae et al. 2023; Ruzza et al. 2024) protoplanets and their hosting disc (e.g., Zhang et al., 2018). However, despite many attempts using infrared (IR) high-contrast imaging (e.g., Testi et al., 2015; Guidi et al., 2018), photospheric emission of accreting planets was detected only in one case (PDS 70, e.g., Keppler et al. 2018; Haffert et al. 2019). On the other hand, a plethora of alternative mechanisms were put forward to explain the wealth of observed substructures (e.g., Bae et al., 2023), such as (magneto-)hydrodynamic instabilities (Lesur et al. 2023 and references therein), thermal and magnetic winds (Pascucci et al. 2023 and references therein), eccentric modes (sometimes triggered by gravitationally bound companions or flybys, e.g., Ragusa et al. 2017; Cuello et al. 2023), or physical-chemical processes that alter the properties of dust at the snowlines of the major volatile species (e.g., Zhang et al., 2015; Okuzumi et al., 2016).
@@
Determining the properties of disc substructures, such as dust density and size (see below), gap/ring contrast (Huang et al., 2018b), trapping ability (Dullemond et al., 2018; Rosotti et al., 2020; Doi & Kataoka, 2023) is of paramount importance not only to discriminate between their possible origins (Bae et al., 2023), but also to assess their potential to form (potentially second generation, see the discussion of Bae et al. 2023) planet(esimal)s, by streaming instability (Youdin & Goodman, 2005; Johansen et al., 2007, see Carrera et al. 2021, 2022; Xu & Bai 2022 for the specific case of pressure bumps), or gravitational instability, (Goldreich & Ward, 1973), and the subsequently growth by pebble accretion (Lau et al., 2022; Jiang & Ormel, 2023).
@@
The substructured discs observed with ALMA at high angular resolution (e.g., Andrews et al., 2018) showed clear indications of moderate-to-high absorption optical depths in the inner regions and at the position of the bright rings (Huang et al., 2018b; Dullemond et al., 2018), considered to be an indication of unresolved optically thick substructures (Dullemond et al., 2018), self-regulating planetesimal formation (Stammler et al., 2019), or high extinction in the presence of self-scattering by highly reflective grains (e.g., Zhu et al., 2019, although Isella et al. 2018 and Guzmán et al. 2018 showed that dust rings do not completely suppress line emission). The spectral index radial profile modulations at the position of gaps and rings inferred from high angular resolution multi-frequency continuum observations (e.g., ALMA Partnership et al., 2015; Tsukagoshi et al., 2016; Macías et al., 2019; Huang et al., 2020a) and the anomalously low ($\alpha<2$) spectral indices in the innermost regions of a few sources (e.g., Huang et al., 2018a; Dent et al., 2019; Paneque-Carreño et al., 2021; Ribas et al., 2023; Houge et al., 2024), most often attributed to dust self-scattering (Liu et al., 2019), go well in line with this picture of moderate-to-high optical depths in the inner disc and the bright rings. These clues suggest that the classically adopted assumption of optically thin continuum emission is rather inaccurate when analysing (sub-)mm observations, and leads to overestimating the grain size and underestimating the dust density (e.g., Zhu et al., 2019; Carrasco-González et al., 2019).
@@
Broadband high angular resolution multi-frequency continuum observations with ALMA (Macías et al., 2021; Sierra et al., 2021; Maucó et al., 2021; Ueda et al., 2022; Ohashi et al., 2023; Carvalho et al., 2024; Sierra et al., 2025), in some cases complemented with Very Large Array (VLA) data (Carrasco-González et al., 2019; Guidi et al., 2022; Zhang et al., 2023; Sierra et al., 2024; Guerra-Alvarado et al., 2024), allowed breaking the degeneracies between intrinsic dust properties and the wavelength dependence of the optical depth, opening up the possibility of thoroughly characterising the temperature, density, and size of solids in gaps and rings. Not surprisingly, most of these works obtained rather different constraints. While partly due to the intrinsic diversity of the analysed sources, the different modelling hypotheses (e.g., the adopted dust composition and porosity, or the fitting methods) might have impacted their results substantially. In this paper, we combine ALMA and VLA high angular resolution and sensitivity (sub-)mm continuum observations and cm-wavelength photometry of CI Tau (whose stellar parameters are summarised in footnote 1), to provide constraints on the properties of solids and simultaneously explore different assumptions on dust composition and porosity.
@@ 1.1 The case of CI Tau
Table 1: Summary of CI Tau’s stellar properties.
(pc) (K) ($L_{\odot}$) ($M_{\odot}$) (yr) ($M_{\odot}\,{\rm yr}^{-1}$)
(1) (2) (3) (4) (5) (6) (7) (8) (9) (10) (11)
CI Tau J04335200+2250301 Taurus $159.6\pm 0.7$ K3 $4559\pm 114$ $0.02\pm 0.20$ $0.11\pm 0.15$ $>3\times 10^{6}$ $-7.28\pm 0.40$ 1, 2
@@
Since the in-situ formation of this candidate planet was considered unlikely (see Rosotti et al. 2017 and references therein), planet migration and interactions with or scattering by an unseen companion were put forward as possible explanations for the origin of this hot Jupiter. On the one hand, while early simulations suggested that the planet eccentricity can be pumped up while migrating in a low-mass disc over long timescales (Rosotti et al., 2017; Ragusa et al., 2018), it was recently shown that type II migration preferably leads to the pile-up of planets between 1 and $10\,{\rm au}$, precluding the formation of hot Jupiters by migration in a (non-self-gravitating) disc (Scardoni et al., 2022). On the other hand, high-contrast IR angular differential imaging ruled out the presence of companions more massive than 2 to $4\,M_{\rm Jup}$ external to $30\,{\rm au}$ (Shimizu et al., 2023), in line with the prediction that eccentricity-driving lower-mass companions are ejected or accreted after dynamical interactions (Rosotti et al., 2017).
Most recently, (sub-)au resolution VLTI/GRAVITY observations (Gravity Collaboration et al., 2023) revealed that CI Tau’s inner disc emission is non-axisymmetric, in line with the predictions of planet-disc interaction simulations (e.g., Muley & Dong, 2021), and misaligned by $\approx 70^{\circ}$ with the outer disc (cf. Clarke et al. 2018), potentially due to the gravitational torque induced by a massive close-in companion. Moreover, the inner dust disc rim is located two to four times further out than the expected position of the sublimation radius, as expected from the presence of an inner companion carving a gap in the disc (Muley & Dong, 2021). This hypothesis was further supported by the shape of the CO $\nu=1-0$ ro-vibrational transitions targeted by IRTF/iSHELL and Keck/NIRSPEC (Kozdon et al., 2023). These lines show a blueshifted core and redshifted wings, that can be explained by two eccentric disc components with oppositely aligned arguments of periapsis, separated by a massive eccentric planet, albeit with much lower eccentricity ($\approx 0.05$) than estimated by Johns-Krull et al. (2016).
@@
Being bright and large at (sub-)mm wavelengths, CI Tau was extensively targeted in the pre-ALMA era, both in single dish (e.g., Beckwith et al., 1990; Beckwith & Sargent, 1991; Andrews & Williams, 2005) and interferometric observations (e.g., Dutrey et al., 1996; Andrews & Williams, 2007; Andrews et al., 2013; Kwon et al., 2015). Early multi-wavelength analyses measured a mm spectral index of $\alpha_{\rm 1.3-2.7mm}=2.5\pm 0.2$ (Dutrey et al., 1996; Ricci et al., 2010b; Guilloteau et al., 2011) that steepens to $\alpha_{\rm 1.3-7.0mm}=3.0\pm 0.3$ at longer wavelengths (Rodmann et al., 2006), and a $\beta^{\rm abs}_{\rm 1.3-2.7mm}$ profile radially increasing between $\lesssim 0.5$ and $\gtrsim 1.0$ (see Figure 10 and 11 of Guilloteau et al. 2011).
@@
As for the lines, although both ${}^{12}{\rm CO}\ J=2-1$ and $J=3-2$ emission is heavily absorbed towards CI Tau (Rosotti et al., 2021; Semenov et al., 2024), a kinematic planetary signature external to the dusty disc was proposed in one of the non-absorbed channels (Rosotti et al., 2021). Instead, ${}^{13}{\rm CO}\ J=2-1$ and $J=3-2$ emission is not absorbed. When observed at high-angular resolution, the latter shows a low-contrast gap co-located with the $\approx 60\,{\rm au}$ continuum ring, attributed by Rosotti et al. (2021) to shadowing by the inner disc. A shallow dip in the ${}^{12}{\rm CO}\ J=3-2$ emission height at the same position of the $\approx 50\,{\rm au}$ gap, and a change of slope co-located with the $\approx 100\,{\rm au}$ ring, were detected by Law et al. (2022). Similarly, the ${}^{12}{\rm CO}$ temperature profile shows two dips at $\approx 70$ and $\approx 120\,{\rm au}$ and a bump at $\approx 90\,{\rm au}$ co-located with similar features in the continuum and ${}^{13}{\rm CO}\ J=3-2$ emission (Law et al., 2022). The ${\rm CS}\ J=7-6$ emission is ring-like, while the $J=5-4$ one is smooth, most likely due to chemical effects (e.g., Le Gal et al., 2019; Rosotti et al., 2021). CCH $N=3-2$ and C18O $J=2-1$ emission are relatively bright and centrally peaked (Bergner et al., 2019; Semenov et al., 2024).

### v2

@@
11institutetext: Max Planck Institute for Astronomy, Königstuhl 17, 69117 Heidelberg, Germany 22institutetext: Institute of Astronomy, University of Cambridge, Madingley Road, Cambridge CB3 OHA, UK 33institutetext: Università degli Studi di Milano, Via Giovanni Celoria 16, 20133 Milano, Italy 44institutetext: European Southern Observatory, Karl-Schwarzschild-Str. 2, 85748 Garching bei München, Germany 55institutetext: Departamento de Astronomía, Universidad de Chile, Camino El Observatorio $15$, Las Condes, Santiago, Chile 66institutetext: Institute for Astronomy, University of Hawaii, Honolulu, HI 96822, USA 77institutetext: INAF – Osservatorio Astrofisico di Arcetri, L.go E. Fermi 5, 50125 Firenze, Italy 88institutetext: Astrophysics Group, Department of Physics, Imperial College London, Prince Consort Rd, London SW7 2A2, UK 99institutetext: School of Physics and Astronomy, University of Leeds, Leeds, LS2 9JT, UK 1010institutetext: Leiden Observatory, Leiden University, P.O. Box 9513, 2300 RA Leiden, The Netherlands 1111institutetext: School of Physics & Astronomy, University of Leicester, University Road, Leicester LE1 7RH, UK 1212institutetext: Dipartimento di Fisica e Astronomia, Università di Bologna, Via Gobetti 93/2, I-40122 Bologna, Italy
@@
Preference for submillimetre-sized low-porosity amorphous carbon grains
@@
We present high angular resolution ($50\,{\rm mas}$) and sensitivity Atacama Large Millimeter/submillimeter Array (ALMA) Band 3 ($3.1\,{\rm mm}$) and Very Large Array (VLA) Ka band ($9.1\,{\rm mm}$) observations of the multi-ringed disc around the $3\,{\rm Myr}$-old solar-mass star CI Tau. These new data were combined with similar-resolution archival ALMA Band 7 ($0.9\,{\rm mm}$) and 6 ($1.3\,{\rm mm}$) observations and new and archival VLA Q ($7.1\,{\rm mm}$), Ku ($2.0\,{\rm cm}$), X ($3.0\,{\rm cm}$) and C band ($6.0\,{\rm cm}$) photometry to study the properties of dust in this system. At wavelengths $\leq 3.1\,{\rm mm}$, the continuum emission from CI Tau is very extended ($\geq 200\,{\rm au}$) and highly substructured (with three gaps, four rings, and two additional gap-ring pairs identified by non-parametric visibility modelling). In contrast, the VLA Ka band data are dominated by a centrally peaked bright component, only partially ($\leq 50\%$) due to dust emission, surrounded by a marginally detected faint and smooth halo. We fitted the ALMA and VLA Ka band data together, adopting a physical model that accounts for the effects of dust absorption and scattering. For our fiducial dust composition (‘Ricci’ opacities), we retrieved a flat maximum grain size distribution across the disc radius, with $a_{\rm max}=(7.1\pm 0.8)\times 10^{-2}\,{\rm cm}$, that we tentatively attributed to fragmentation of fragile dust or bouncing. We tested, for the first time, the dependence of our results on the adopted dust composition model to assess which mixture can best reproduce the observations. We found that ‘Ricci’ opacities work better than the traditionally adopted ‘DSHARP’ ones, while graphite-rich mixtures perform significantly worse. We also show that for our fiducial composition, the data prefer low porosity ($\leq 70\%$) grains. This is in contrast with recent claims of highly porous aggregates in younger sources, which we tentatively justified by time-dependent compaction at the fragmentation or bouncing barrier. Our results on composition and porosity are in line with constraints from disc population synthesis models and naturally arise from CI Tau’s peculiar spectral behaviour (i.e. the abrupt steepening of its spectral index at wavelengths longer than $3.1\,{\rm mm}$), making this disc a unique target to characterise the properties of disc solids and thus ideal for deeper centimetre-wavelength observations and follow-up dust polarisation studies.
@@ 1 Introduction
Dust is an essential ingredient of every planet-formation model. First and foremost, it is the material planet(esimal)s are made of. By dominating the disc opacity, solids also set the temperature structure and thus the excitation conditions and reaction rates of volatiles (e.g. Gavino et al., 2021; Öberg et al., 2023). In addition, electron-ion recombination and chemical reactions on the surface of grains affect the disc ionisation structure (e.g. Desch & Turner, 2015) and prompt the formation of complex organic molecules (e.g. Öberg & Bergin, 2021). Dust collisional evolution (growth and fragmentation) and transport (vertical settling and radial drift) not only determine the availability of solids to form planetary cores (Brauer et al., 2008; Birnstiel et al., 2010) but also affect the dust-to-gas thermal coupling (Facchini et al., 2017) and the abundance of gas-phase chemical species. As dust migrates radially, the volatiles frozen onto grains undergo thermal desorption upon transition across their snowlines (e.g. Booth et al., 2017; Booth & Ilee, 2019; Krijt et al., 2020; Van Clepper et al., 2022), thus locally enriching the feedstock of exoplanet atmospheres (Madhusudhan, 2019). Therefore, unsurprisingly, determining the physical and chemical properties of protoplanetary disc solids is an essential first step to understanding planet formation (e.g. Miotello et al., 2023; Birnstiel, 2024). For this reason, attempts have been made to infer the density, size, distribution, and temperature of disc solids by modelling the continuum emission from planet-forming discs at multiple wavelengths.
@@
Early works, which either focused on a few targets of interest or surveyed tens of discs in nearby star-formation regions, measured integrated spectral indices, first in the (sub-)millimetre and then at progressively longer wavelengths significantly smaller than in the interstellar medium (ISM). While the contribution of optically thick regions could not be ruled out (Ricci et al., 2012), especially at (sub-)millimetre wavelengths, these results were generally interpreted (Beckwith & Sargent, 1991; Andrews & Williams, 2005; Rodmann et al., 2006), sometimes in combination with information on the (wavelength dependence of the) disc radius (e.g. Testi et al., 2003; Wilner et al., 2005; Ricci et al., 2010a, b; Isella et al., 2010; Banzatti et al., 2011; Guilloteau et al., 2011), as being a consequence of substantial grain growth up to millimetre or centimetre sizes, depending on the adopted dust optical properties (see Testi et al. 2014 and Andrews 2020 for a summary picture). In addition, the first multi-wavelength studies able to resolve a few bright sources from (sub-)millimetre to centimetre wavelengths (Pérez et al., 2012, 2015; Trotta et al., 2013; Tazzari et al., 2016) inferred radially decreasing grain size profiles, in qualitative agreement with the predictions of dust evolution models (Brauer et al., 2008; Birnstiel et al., 2010).
@@
The advent of the Atacama Large Millimeter/submillimeter Array (ALMA) substantially revolutionised this picture. One of ALMA’s major breakthroughs was the almost ubiquitous detection of small scale structures in the continuum emission of (the largest and brightest) protoplanetary discs (Andrews, 2020; Bae et al., 2023). At (sub-)millimetre wavelengths, substructures are most commonly detected in the form of axisymmetric troughs and peaks (colloquially referred to as ‘gaps’ and ‘rings’; e.g. Andrews et al. 2016; Long et al. 2018; Andrews et al. 2018; Huang et al. 2018b) or cavities (e.g. Keppler et al., 2019; Facchini et al., 2020), but spirals (Pérez et al., 2016; Huang et al., 2018c; Kurtovic et al., 2018; Paneque-Carreño et al., 2021) and crescents (e.g. van der Marel et al., 2013; Isella et al., 2018; Pérez et al., 2018; Dong et al., 2018) are sometimes also present. On the one hand, the most tantalising interpretation of these structures is that they trace dynamical interactions between massive (0.1 to $10.0\,M_{\rm Jup}$, Lodato et al. 2019; Bae et al. 2023; Ruzza et al. 2024) protoplanets and their hosting disc (e.g. Zhang et al., 2018). However, despite many attempts using infrared (IR) high-contrast imaging (e.g. Testi et al., 2015; Guidi et al., 2018), photospheric emission of accreting planets was detected only in one case (PDS 70, e.g. Keppler et al. 2018; Haffert et al. 2019). On the other hand, a plethora of alternative mechanisms have been put forward to explain the wealth of observed substructures (e.g. Bae et al., 2023), such as (magneto-)hydrodynamic instabilities (Lesur et al. 2023 and references therein), thermal and magnetic winds (Pascucci et al. 2023 and references therein), eccentric modes (sometimes triggered by gravitationally bound companions or flybys, e.g. Ragusa et al. 2017; Cuello et al. 2023), or physical-chemical processes that alter the properties of dust at the snowlines of the major volatile species (e.g. Zhang et al., 2015; Okuzumi et al., 2016).
@@
Determining the properties of disc substructures, such as dust density and size (see below), gap-to-ring contrast (Huang et al., 2018b), and trapping ability (Dullemond et al., 2018; Rosotti et al., 2020; Doi & Kataoka, 2023), is of paramount importance. This is crucial not only to discriminate between their possible origins (Bae et al., 2023) but also to assess their potential to form (second generation, see the discussion of Bae et al. 2023) planet(esimal)s by streaming instability (Youdin & Goodman, 2005; Johansen et al., 2007, see Carrera et al. 2021, 2022; Xu & Bai 2022 for the specific case of pressure bumps) or gravitational instability (Goldreich & Ward, 1973) and the subsequent growth by pebble accretion (Lau et al., 2022; Jiang & Ormel, 2023).
@@
The substructured discs observed with ALMA at high angular resolution (e.g. Andrews et al., 2018) showed clear indications of moderate-to-high absorption optical depths in the inner regions and at the position of the bright rings (Huang et al., 2018b; Dullemond et al., 2018), which is considered to be an indication of unresolved optically thick substructures (Dullemond et al., 2018), self-regulating planetesimal formation (Stammler et al., 2019), or high extinction in the presence of self-scattering by highly reflective grains (e.g. Zhu et al., 2019, although Isella et al. 2018 and Guzmán et al. 2018 showed that dust rings do not completely suppress line emission). The spectral index radial profile modulations at the position of gaps and rings inferred from high angular resolution multi-frequency continuum observations (e.g. ALMA Partnership et al., 2015; Tsukagoshi et al., 2016; Mac´ıas et al., 2019; Huang et al., 2020a) and the ‘anomalously low’ (i.e. $\alpha<2$) spectral indices in the innermost regions of a few sources (e.g. Huang et al., 2018a; Dent et al., 2019; Paneque-Carreño et al., 2021; Ribas et al., 2023; Houge et al., 2024), which are most often attributed to dust self-scattering (Liu, 2019), go well in line with this picture of moderate-to-high optical depths in the inner disc and the bright rings. These clues suggest that the classically adopted assumption of optically thin continuum emission is rather inaccurate when analysing (sub-)millimetre observations and leads to overestimation of the grain size and underestimation of the dust density (e.g. Zhu et al., 2019; Carrasco-González et al., 2019).
@@
Broadband high angular resolution multi-frequency continuum observations with ALMA (Mac´ıas et al., 2021; Sierra et al., 2021; Maucó et al., 2021; Ueda et al., 2022; Ohashi et al., 2023; Carvalho et al., 2024; Sierra et al., 2025), complemented with Very Large Array (VLA) data in some cases (Carrasco-González et al., 2019; Guidi et al., 2022; Zhang et al., 2023; Sierra et al., 2024; Guerra-Alvarado et al., 2024), allowed the degeneracies between intrinsic dust properties and the wavelength dependence of the optical depth to be broken, thus opening up the possibility of thoroughly characterising the temperature, density, and size of solids in gaps and rings. Not surprisingly, most of these works obtained rather different constraints. While partly due to the intrinsic diversity of the analysed sources, the different modelling hypotheses (e.g. the adopted dust composition and porosity or the fitting methods) might have impacted their results substantially. In this paper, we combine ALMA and VLA high angular resolution and sensitivity (sub-)millimetre continuum observations and centimetre-wavelength photometry of CI Tau (whose stellar parameters are summarised in footnote 20) to provide constraints on the properties of solids and simultaneously explore different assumptions on dust composition and porosity.
@@ 1.1 The case of CI Tau
Since the in situ formation of this candidate planet was considered unlikely (see Rosotti et al. 2017 and references therein), planet migration and interactions with or scattering by an unseen companion were put forward as possible explanations for the origin of this hot Jupiter. On the one hand, while early simulations suggested that the planet eccentricity can be pumped up while migrating in a low-mass disc over long timescales (Rosotti et al., 2017; Ragusa et al., 2018), it was recently shown that type II migration preferably leads to the pile-up of planets between 1 and $10\,{\rm au}$, precluding the formation of hot Jupiters by migration in a (non-self-gravitating) disc (Scardoni et al., 2022). On the other hand, high-contrast IR angular differential imaging ruled out the presence of companions more massive than 2 to $4\,M_{\rm Jup}$ external to $30\,{\rm au}$ (Shimizu et al., 2023), in line with the prediction that eccentricity-driving lower-mass companions are ejected or accreted after dynamical interactions (Rosotti et al., 2017).
Most recently, (sub-)au resolution VLTI/GRAVITY observations (Gravity Collaboration et al., 2023) revealed that CI Tau’s inner disc emission is non-axisymmetric, in line with the predictions of planet-disc interaction simulations (e.g. Muley & Dong, 2021), and misaligned by $\approx 70^{\circ}$ with the outer disc (see Clarke et al. 2018), potentially due to the gravitational torque induced by a massive close-in companion. Moreover, the inner dust disc rim is located two to four times further out than the expected position of the sublimation radius, as expected from the presence of an inner companion carving a gap in the disc (Muley & Dong, 2021). This hypothesis was further supported by the shape of the CO $\nu=1-0$ ro-vibrational transitions targeted by IRTF/iSHELL and Keck/NIRSPEC (Kozdon et al., 2023). These lines show a blueshifted core and redshifted wings, that can be explained by two eccentric disc components with oppositely aligned arguments of periapsis, separated by a massive eccentric planet, albeit with much lower eccentricity ($\approx 0.05$) than estimated by Johns-Krull et al. (2016).
@@
Being bright and large at (sub-)millimetre wavelengths, CI Tau was extensively targeted in the pre-ALMA era, both in single dish (e.g. Beckwith et al., 1990; Beckwith & Sargent, 1991; Andrews & Williams, 2005) and interferometric observations (e.g. Dutrey et al., 1996; Andrews & Williams, 2007; Andrews et al., 2013; Kwon et al., 2015). Early multi-wavelength analyses measured a millimetre spectral index of $\alpha_{\rm 1.3-2.7mm}=2.5\pm 0.2$ (Dutrey et al., 1996; Ricci et al., 2010b; Guilloteau et al., 2011) that steepens to $\alpha_{\rm 1.3-7.0mm}=3.0\pm 0.3$ at longer wavelengths (Rodmann et al., 2006), and a $\beta^{\rm abs}_{\rm 1.3-2.7mm}$ profile radially increasing between $\lesssim 0.5$ and $\gtrsim 1.0$ (see Figure 10 and 11 of Guilloteau et al. 2011).
@@
As for the lines, although both $^{12}{\rm CO}\ J=2-1$ and J=32J=3-2 emission is heavily absorbed towards CI Tau (Rosotti et al., 2021; Semenov et al., 2024), a kinematic planetary signature external to the dusty disc was proposed in one of the non-absorbed channels (Rosotti et al., 2021). Instead, $^{13}{\rm CO}\ J=2-1$ and J=32J=3-2 emission is not absorbed. When observed at high-angular resolution, the latter shows a low-contrast gap co-located with the $\approx 60\,{\rm au}$ continuum ring, attributed by Rosotti et al. (2021) to shadowing by the inner disc. A shallow dip in the $^{12}{\rm CO}\ J=3-2$ emission height at the same position of the $\approx 50\,{\rm au}$ gap, and a change of slope co-located with the $\approx 100\,{\rm au}$ ring, were detected by Law et al. (2022). Similarly, the $^{12}{\rm CO}$ temperature profile shows two dips at $\approx 70$ and $\approx 120\,{\rm au}$ and a bump at $\approx 90\,{\rm au}$ co-located with similar features in the continuum and $^{13}{\rm CO}\ J=3-2$ emission (Law et al., 2022). The ${\rm CS}\ J=7-6$ emission is ring-like, while the J=54J=5-4 one is smooth, most likely due to chemical effects (e.g. Le Gal et al., 2019; Rosotti et al., 2021). CCH N=32N=3-2 and C18O J=21J=2-1 emission are relatively bright and centrally peaked (Bergner et al., 2019; Semenov et al., 2024).